U = 4x11/2 + 2x2. p1 = 12 and p2 = 8. Income (m) is 173.16. The utility optimizing quantity of x2 is
Answer #1

At optimal; MRS = p1/p2

MRS = MUx1/MUx2

MUx1 = 4 * 1/2 x1(-1/2) = 2/(x1)1/2

MUx2 = 2

So, (2/(x1)1/2)/2 = 12/8

1/(x1)1/2 = 12/8

(x1)1/2 = 8/12 = 2/3 = 0.67

x1 = 0.45

Budget line: p1x1 + p2x2 = m

12x1 + 8x2 = 173.16

12(0.45) + 8x2 = 173.16

8x2 = 173.16 - 5.4

x2 = 20.97